Help please :) I really don't want the answer....please just help me w/ the proper steps to solve.
A child in an airport is able to cover meters in minutes running at a steady speed down a moving sidewalk in the direction of the sidewalk's motion. Running at the same speed in the direction opposite to the sidewalk's movement, the child is able to cover meters in minutes. What is the child's running speed on a still sidewalk, and what is the speed of the moving sidewalk?

Answer by venugopalramana(3286)

Help please :) I really don't want the answer....please just help me w/ the proper steps to solve.
A child in an airport is able to cover meters in minutes running at a steady speed down a moving sidewalk in the direction of the sidewalk's motion. Running at the same speed in the direction opposite to the sidewalk's movement, the child is able to cover meters in minutes. What is the child's running speed on a still sidewalk, and what is the speed of the moving sidewalk?
VERY GOOD TO SEE YOUR KEEN DESIRE TO LEEARN..IN THESE TYPES OF TIME AND DISTANCE PROBLEMS,WE USE NORMAL FORMULAE WITH SLIGHT MODIFICATION
STANDARD FORMULA FOR TIME AND DISTANCE IS
DISTANCE TRAVELLED = D = SPEED * TIME =S*T
REVISED FORMULA WE USE IN GENERAL CASES IS
DISTANCE OF SEPERATION FROM INITIAL STAGE TO FINAL STAGE=D2-D1=D
TIME = TIME AT END - TIME AT THE BEGINING =T2-T1=T
RELATIVE SPEED = SUM OF SPEEDS WHEN THE SPEEDS OF 2 MOVING OBJECTS AID EACH OTHER ....... = S1+S2=S
RELATIVE SPEED = DIFFERENCE IN SPEEDS WHEN THE SPEEDS OF 2 MOVING OBJECTS OPPOSE EACH OTHER ....... = S1-S2=S
LET US EXPAND A BIT ON THIS RELATIVE SPEED..SUPPOSE YOU WALK AT 3 MPH IN A MOVING TRAIN IN THE SAME DIRECTION AS TRAIN IS MOVING,THEN WITH RESPECT TO THE TRAIN YOU ARE ONLY WALKING AT 3 MPH, BUT W.R.T. A STATIONARY GROUND ,YOUR RELATIVE SPEED WILL BE 3 MPH+SPEED OF TRAIN SAY 50 MPH =53 MPH.
SIMILARLY WHEN YOU ARE WALKING IN THE OPPOSITE DIRECTION TO THAT OF TRAIN ,YOUR RELATIVE SPEED W.R.T.TO THE GROUND =50-3=47 MPH.
SAME THING APPLIES WHEN YOU ARE SWIMMING IBN A CURRENT OF WATER OR WALKING ON A MOVING SIDE WALK OR A PLANE GOES AGAINST THE WIND OR PROWIND ETC...
IN YOUR PRESENT PROBLEM YOU CAN TAKE
LET SPEED OF CHILD ON STILL SIDE WALK =S1 AND
SPEED OF MOVING SIDE WALK =S2
SO WHEN THEY ARE GOING IN THE SAME DIRECTION RELATIVE SPEED W.R.T GROUND =S1+S2
& WHEN THEY ARE GOING IN OPPOSITE DIRECTIONS RELATIVE SPEED W.R.T GROUND =S1-S2
HENCE TIME TAKEN WILL BE D/(S1+S2) IN ONE CASE AND D/(S1-S2) IN THE OPPOSITE CASE.HOPE YOU CAN SOLVE YOUR PROBLEM NOW.